OpenStudy (acacia):

Help help help!! Find the range of the logarithmic function f(x) = log4 (x – 3)

OpenStudy (anonymous):

Remember x > 0 Domain of log function = \[(0,\infty) \] Range of log functions is \[(-\infty,\infty)\] So we must find where the x value in your equation is greater than 0 So solve for x-3 > 0